OpenStudy (anonymous):

using mileage distance: how to find the distance between (39.970, -83.001) from location (39.579, -82.373)

OpenStudy (anonymous):

\[\sqrt{(x2-x1)^2+(y2-y1)^2}\] You have to use the distance formula for this problem. \[\sqrt{(39.579-39.970)^2+(-82.373-(-83.001))^2}\]

OpenStudy (anonymous):

All you have to do is solve that equation and you'll get your answer :)

OpenStudy (anonymous):

Okay so I got this can someone check my work? sqrt{(69.1(39.579-39.970))^2+(53(-82.373-(-83.110))^2 sqrt{(69.1*(-0.391)^2+(53*(0.628)^2 sqrt{(69.1*(-0.153)+(53*.394) 10.572+20.882=31.453 =31.454 so would be 5.61 miles from one town to another